Based on economists’ forecasts and analysis, 1-year Treasury bill rates and liquidity premiums for the next four years are expected to be as follows:

R1 = 0.90 % E(2r1) = 2.05 % L2 = 0.09 % E(3r1) = 2.15 % L3 = 0.12 % E(4r1) = 2.45 % L4 = 0.14 %

Using the liquidity premium theory, determine the current (long-term) rates; for FOUR years. (Do not round intermediate calculations. Round your answers to 2 decimal places.)
Answer #1

R1 = 0.90%

R2 = [(1 + R1) * {1 + E(2r1)}]1/2 - 1 + L2

= [(1 + 0.009) * (1 + 0.0205)]1/2 - 1 + 0.0009

= [1.0297]1/2 - 0.9991 = 1.0147 - 0.9991 = 0.0156, or 1.56%

R3 = [(1 + R1) * {1 + E(2r1)} * {1 + E(3r1)}]1/3 - 1 + L3

= [(1 + 0.009) * (1 + 0.0205) * (1 + 0.0215)]1/3 - 1 + 0.0012

= [1.0518]1/3 - 0.9988 = 1.0170 - 0.9988 = 0.0182, or 1.82%

R4 = [(1 + R1) * {1 + E(2r1)} * {1 + E(3r1)} * {1 + E(3r1)}]1/4 - 1 + L4

= [(1 + 0.009) * (1 + 0.0205) * (1 + 0.0215) * (1 + 0.0245)]1/4 - 1 + 0.0014

= [1.0776]1/4 - 0.9986 = 1.0189 - 0.9986 = 0.0203, or 2.03%